Consider the following table for a? household's consumption expenditures and disposable income.

To the nearest? dollar, compute desired saving at each level of disposable income. ?(Enter your responses as whole numbers and include a minus sign where? appropriate.)

Income

Consumption

Savings

?$0

?$100

?$

?$100

?$150

?$

?$200

?$200

?$

?$300

?$250

?$

?$400

?$300

?$

?$500

?$350

?$

The marginal propensity to save is ____.

?(Enter your response rounded to two decimal? places.)

The marginal propensity to save ____ along a given saving schedule.

The marginal propensity to save plus the marginal propensity to consume must equal ____.

?(Enter your response rounded to two decimal? places.)
Answer #1

Savings=Income-Consumption

Income Consumption Saving
0 100 -100
100 150 -50
200 200 0
300 250 50
400 300 100
500 350 150

The marginal propensity to save is 0.50.
MPS=Change in savings/Change in income
MPS=(50-0)/(300-200)
MPS=50/100
MPS=0.50

The marginal propensity to save remains constant along a given saving schedule.
MPS=(100-50)/(400-300)
MPS=50/100
MPS=0.50
MPS between income 300 and 200 is 0.5 and between income 400 and 300 is also 0.5.

The marginal propensity to save plus the marginal propensity to consume must equal 1.
MPC+MPS=1